(D^3 - 27) = e^3x + e^-3x
1
Expert's answer
2020-12-11T09:47:04-0500

"y^{(3)}-27y=e^{3x}+e^{-3x}"

First we solve the homogenous equation "y^{(3)}-27y=0"

The characterisic equation is "\lambda^3-27=0", which has roots "\lambda_1=3, \lambda_2=3e^\frac{2i\pi}{3}=-\frac{3}{2}+\frac{3\sqrt{3}}{2}i, \lambda_3=3e^\frac{-2i\pi}{3}=-\frac{3}{2}-\frac{3\sqrt{3}}{2}i"

So the solution of "y^{(3)}-27y=0" is "y=Ae^{3x}+Be^{-\frac{3}{2}x}\cos\frac{3\sqrt{3}}{2}x+Ce^{-\frac{3}{2}x}\sin\frac{3\sqrt{3}}{2}x"

Now we find partial solution of "y^{(3)}-27y=e^{3x}+e^{-3x}" in form "y=Vxe^{3x}+We^{-3x}"

By the General Leibniz rule we have "(xe^{3x})^{(3)}=x^{(3)}e^{3x}+3x''(e^{3x})'+3x'(e^{3x})''+x(e^{3x})^{(3)}="

"=0+0+3\cdot 9e^{3x}+27xe^{3x}=27e^{3x}+27xe^{3x}"

So "y^{(3)}-27y=27Ve^{3x}+27Vxe^{3x}-27We^{-3x}-"

"-27Vxe^{3x}-27We^{-3x}=27Ve^{3x}-54We^{-3x}"

Then "27V=-54W=1", that is "V=\frac{1}{27},W=-\frac{1}{54}"

We obtain the general solution "y=Ae^{3x}+Be^{-\frac{3}{2}x}\cos\frac{3\sqrt{3}}{2}x+Ce^{-\frac{3}{2}x}\sin\frac{3\sqrt{3}}{2}x+"

"+\frac{1}{27}xe^{3x}-\frac{1}{54}e^{-3x}"

Answer: "Ae^{3x}+Be^{-\frac{3}{2}x}\cos\frac{3\sqrt{3}}{2}x+Ce^{-\frac{3}{2}x}\sin\frac{3\sqrt{3}}{2}x+"

"+\frac{1}{27}xe^{3x}-\frac{1}{54}e^{-3x}"